OpenStudy (anonymous):

Write an equation for a line perpendicular to y=3/4x-4 and goes through the point (1,-2)

OpenStudy (anonymous):

Wells the slope is reciprocal of 3/4 which is -4/3 as the slope. We then plug in values for the left and right side resulting in -2=-4/3+a which is basically a=-2/3 thus the equation is y=-4/3x-2/3

OpenStudy (anonymous):

Reward me please.